CUSO4 + Zn ———> Cu + ZnSO4

how many grams of copper could be obtained by the reaction of 32 g of cuso4 with an excess of zinc?
1
Expert's answer
2018-09-06T04:48:20-0400
At first we must calculate the mole of CuSO4.
n(CuSO4)=32/160= 0.2 moles of CuSO4.
From reaction we can see that one mole of CuSO4 gives one mole of Cu. It means that 0.2 moles of CuSO4 will give 0.2 mole of Cu.
Then we can calculate the mass of Cu.
m(Cu)=0.2*64=12.8 grams of Cu.